# Practical Dirac Majorana confusion theorem: Issues and Applicability

### v1

@@ 5 Discussions on concepts usually accompanying explanations of pDMCT
5.1 Is there any one-to-one correspondence between Dirac and Majorana neutrinos in the massless limit, $m_{\nu}\to 0$
@@
The issue of one-to-one correspondence between Dirac and Majorana neutrinos in the massless limit $m_{\nu}\to 0$ can also be analyzed from the context of specific processes and observables. As mentioned in Sec. 2.3 if one considers neutral current mediated processes such as $Z\to\nu\,\overline{\nu}$ (and include summation over neutrino spins while evaluating amplitude squares) within the SM, the direct and exchange terms in amplitude square become equal and the difference between Dirac and Majorana neutrinos becomes proportional to $m_{\nu}^{2}$, which vanishes if we were to simply apply the limit $m_{\nu}\to 0$ at the end. [^11] As explained before in Appendix B, we can not start from massless neutrinos as then the neutrinos can not have Majorana nature. At the end of full calculation, one can certainly apply the limit $m_{\nu}\to 0$ which amounts to neglecting $m_{\nu}$ dependent terms due to their tininess. However, when one considers processes that are not facilitated by the SM neutral current interactions, the direct and exchange terms can have non-zero difference even when we neglect $m_{\nu}$ dependent terms (or equivalently put $m_{\nu}\to 0$). In such a case some specific observable might be able to probe these important non-zero differences. In these instances, in the context of a specific process and observable, there is indeed no one-to-one correspondence between the Dirac and Majorana neutrino scenarios. However, in all cases with the SM only interactions if the observable includes full phase space integration over the neutrino and antineutrino, we do find that the direct and exchange terms have equal contribution after integration (see Eq. (8)) which amounts to no observable difference between the two scenarios in the limit $m_{\nu}\to 0$.
@@
As explained in Appendix B, it is a mathematical impossibility to preserve Majorana nature of a fermion when its mass is zero. In fact due to Lorentz invariance and conservation of chirality for massless fermions, such chiral fermions have a distinct nomenclature as being Weyl fermions. To describe Weyl fermions it is sufficient to use 2-component complex spinors instead of 4-component complex spinors. Nevertheless, the 2-component Weyl spinors can be used to construct 4-component complex Dirac spinors as well as 4-component Majorana spinors (that are real in the Majorana basis). Both the Dirac and Majorana spinors have both left- and right-chiral components. When one takes the massless limit or when one considers ultra-relativistic fermions one finds that these constructs of Dirac or Majorana spinors prefer specific chirality states. Nevertheless, as long as the mass of the fermion is non-zero both the chiral states are present. However, once the fermion is massless, it becomes fully chiral and it can not have Majorana nature at all. The Dirac nature (meaning its particle and antiparticle states are distinct and distinguishable) survives the massless limit. Therefore, although both Dirac and Majorana 4-component spinors get reduced to 2-component Weyl spinors in the massless limit, the Weyl spinors only show Dirac nature and the Majorana nature is completely lost. This implies there is really no one-to-one correspondence between Dirac and Majorana nature of neutrinos in the massless limit.
@@
5.2 Should the amplitude be anti-symmetrized for pair of Majorana neutrinos of the same flavor with $m_{\nu}>0$
@@
If the 4-momenta (and spins) of $\nu_{\ell}$, $\overline{\nu}_{\ell}$ be denoted by $p_{1}$, $p_{2}$ (and $s_{1}$, $s_{2}$) respectively, then the transition amplitude for Dirac case can be symbolically be written as,
@@
In the calculation of the observable for the specific process in both Dirac and Majorana cases, one takes the square of the amplitude, does the usual trace calculations by summing over the final spins and averages over the initial spins, except when one is interested in an observable that depends on neutrino spins which is practically impossible to do experimentally for sub-eV active neutrinos. Thus the spin information gets wiped out via the spin summation. In context of the SM, we know that the $V-A$ nature of weak interaction ensures that we always get a left-chiral neutrino and a right-chiral antineutrino. However, despite being produced in specific chiral states, their chirality is not conserved due to non-zero mass, following Eq. (B). Due to non-zero mass, chirality is not same as helicity which is the projection of spin along the direction of motion. Thus, left and right helical massive neutrinos get produced from the SM weak interaction. Thus we consider all spin possibilities of the massive eigenstates in our calculation. Any issues related to relativistic or non-relativistic kinematics are automatically taken care of by the field theoretic calculations for amplitude square with summation over final spins and average over initial spins.
@@
In the Majorana case amplitude square, it thus becomes clear that there will be interference terms that require helicity flip and such terms would be proportional to $m_{\nu}^{2}$. If the SM interactions are only taken into account, then all the interference terms always involve helicity flips. It should be noted that it is by antisymmetrization alone do we get axial-vector current interaction for Majorana neutrinos in $Z\to\nu\,\overline{\nu}$ as opposed to vector minus axial-vector current interaction for Dirac neutrinos.
@@ Acknowledgements
First of all, I would like to thank my previous collaborators Dibyakrupa Sahoo, Janusz Rosiek and M.V.N. Murthy for their helpful discussions and comments. I would also like to thank K. Hagiwara, S. Parke, J. Kopp, G. Lopez-Castro, Z. Xing, A. Smirnov, E. Akhmedov and M. Misiak for their valuable comments and criticism. The work of CSK is supported by NRF of Korea (NRF-2021R1A4A2001897 and NRF-2022R1I1A1A01055643).
@@ A.1 Important issues to address when coming to back-to-back kinematics from general kinematics
The usual description of general kinematics for the 4-body final state has two decay planes, with an angle $\phi$ between them. For back-to-back kinematics the two decay planes coincide to form a single decay plane. So for back-to-back kinematics $\phi=0$. No integration over $\phi$ is involved to arrive at the final observables in back-to-back kinematics.
@@
Out of these, the first and second conditions are primary ones, while the rest arise as a consequence of the first two criteria. However, none of these quantities are physically observable. The first two conditions also imply that $E_{+}=E_{-}=E_{\mu}\equiv\frac{m_{B}}{2}-E_{\nu}$ which is experimentally observable. Any error, on the muon energy measurement would imply that the muon energy distribution as shown in Fig. 5(c) of Ref. Kim:2021dyj would involve energy bins with bin width corresponding to the experimental error within which equality of $E_{+}$ and $E_{-}$ is satisfied. There would also possibly be extremely slight deviation from the $180^{\circ}$ angle between the two final muons, in an experimental back-to-back realization. This can lead to small deviation of $\Theta$ from $\pi$, say $\Theta=\pi\pm\Delta\Theta$. This implies $\cos\Theta=-\cos\Delta\Theta\approx-1+\frac{\Delta\Theta^{2}}{2}$, so that the error in $\cos\Theta$ measurement is $\Delta\Theta^{2}/2$ and it gets multiplied to the back-to-back muon energy distribution of Eq. (50) of Ref. Kim:2021dyj to give the amount of smearing one can expect from the measurement. Thus, a real experimental realization of the back-to-back kinematics would lead to a muon energy distribution curve similar to Fig. 5(c) of Ref. Kim:2021dyj but it would be a histogram plot with bin size determined by the muon energy resolution and there will be some vertical smearing arising from the slight deviation from $180^{\circ}$ angle requirement. However, the difference between Dirac and Majorana neutrinos should not get washed away as a result of such a tiny smearing from experimental measurements.

### v4

@@ 5 Discussions on concepts usually accompanying explanations of pDMCT
5.1 Is there any one-to-one correspondence between Dirac and Majorana neutrinos in the massless limit, $\boldsymbol{m_{\nu}\to 0}$?
@@
The issue of one-to-one correspondence between Dirac and Majorana neutrinos in the massless limit $m_{\nu}\to 0$ can be analyzed from the context of specific processes and observables. As mentioned in Sec. 2.3 if one considers neutral current mediated processes such as $Z\to\nu\,\overline{\nu}$ (and include summation over neutrino spins while evaluating amplitude squares) within the SM, the direct and exchange terms in amplitude square become equal and the difference between Dirac and Majorana neutrinos becomes proportional to $m_{\nu}^{2}$, which vanishes if we were to simply apply the limit $m_{\nu}\to 0$ at the end. [^11] As shown in Appendix B, we can not start from massless neutrinos as then the neutrinos can not have Majorana nature. At the end of full calculation, one can certainly apply the limit $m_{\nu}\to 0$ which amounts to neglecting $m_{\nu}$ dependent terms due to their tininess. However, when one considers processes that are not facilitated by the SM neutral current interactions, the direct and exchange terms can have non-zero difference even when we neglect $m_{\nu}$ dependent terms (or equivalently put $m_{\nu}\to 0$). In such a case some specific observable might be able to probe these important non-zero differences. In these instances, in the context of a specific process and observable, there is indeed no one-to-one correspondence between the Dirac and Majorana neutrino scenarios. However, in all cases with the SM only interactions if the observable includes full phase space integration over the neutrino and antineutrino, we do find that the direct and exchange terms have equal contribution after integration (see Eq. (8)) which amounts to no observable difference between the two scenarios in the limit $m_{\nu}\to 0$.
@@
As explained in Appendix B, it is a mathematical impossibility to preserve Majorana nature of a fermion when its mass becomes zero. In fact due to Lorentz invariance and conservation of chirality for massless fermions, such chiral fermions have a distinct nomenclature as being Weyl fermions. To describe Weyl fermions it is sufficient to use 2-component complex spinors instead of 4-component complex spinors. Nevertheless, the 2-component Weyl spinors can be used to construct 4-component complex Dirac spinors as well as 4-component Majorana spinors (that are real in the Majorana basis). Both the Dirac and Majorana spinors have both left- and right-chiral components. When one takes the massless limit or when one considers ultra-relativistic fermions one finds that these constructs of Dirac or Majorana spinors prefer specific chirality states. Nevertheless, as long as the mass of the fermion is non-zero both the chiral states are present. However, once the fermion is massless, it becomes fully chiral and it can not have Majorana nature at all. The Dirac nature (meaning its particle and antiparticle states are distinct and distinguishable) survives the massless limit. Therefore, although both Dirac and Majorana 4-component spinors get reduced to 2-component Weyl spinors in the massless limit, the Weyl spinors only show Dirac nature and the Majorana nature is completely lost. This implies there is really no one-to-one correspondence between Dirac and Majorana nature of neutrinos in the massless limit.
@@
5.2 Should the amplitude be anti-symmetrized for pair of Majorana neutrinos of the same flavor with $\boldsymbol{m_{\nu}>0}$?
@@
If the 4-momenta (and spins) of $\nu_{\ell}$, $\overline{\nu}_{\ell}$ be denoted by $p_{1}$, $p_{2}$ (and $s_{1}$, $s_{2}$) respectively, then the transition amplitude for Dirac case can be symbolically written as,
@@
In the calculation of the observable for the specific process in both Dirac and Majorana cases, one takes the square of the amplitude, does the usual trace calculations by summing over the final spins and averages over the initial spins, except when one is interested in an observable that depends on neutrino spins which is practically impossible to do experimentally for sub-eV active neutrinos. Thus the spin information gets wiped out via the spin summation. In context of the SM, we know that the $V-A$ nature of weak interaction ensures that we always get a left-chiral neutrino and a right-chiral antineutrino. However, despite being produced in specific chiral states, their chirality is not conserved due to non-zero mass, following Eq. (38). Due to non-zero mass, chirality is not same as helicity which is the projection of spin along the direction of motion. Thus, left and right helical massive neutrinos get produced from the SM weak interaction. Thus we consider all spin possibilities of the massive eigenstates in our calculation. Any issues related to relativistic or non-relativistic kinematics are automatically taken care of by the field theoretic calculations for amplitude square with summation over final spins and average over initial spins.
@@
In the Majorana case amplitude square, it thus becomes clear that there will be interference terms that require helicity flip and such terms would be proportional to $m_{\nu}^{2}$. If the SM interactions are only taken into account, then all the interference terms always involve helicity flips.
@@ 6 Conclusions
As a final conclusion, I would like to stress that quantum statistical effects can be used to pin down the nature of the neutrino – whether it is Majorana or Dirac type. In the context of processes allowed in the Standard Model, this follows when
  1. 1.
    the direct and exchange terms in amplitude square are non-trivially different (see Section 2.2), and
  2. 2.
    the observable defined in terms of neutrino momenta is accessible even when the neutrinos are not directly observed in the final state (see Section 4.2).
@@ Acknowledgements
First of all, I would like to thank my previous collaborators Dibyakrupa Sahoo, Janusz Rosiek and M.V.N. Murthy for their helpful discussions and comments. I would also like to thank K. Hagiwara, S. Parke, J. Kopp, G. Lopez-Castro, Z. Xing, A. Smirnov, E. Akhmedov and M. Misiak for their valuable comments and criticism. The work of CSK is supported by NRF of Korea (NRF-2021R1A4A2001897 and
NRF-2022R1I1A1A01055643).
@@ A.1 Important issues to address when coming to back-to-back kinematics from general kinematics
The usual description of general kinematics for the 4-body final state has two decay planes, with an angle $\phi$ between them. For back-to-back kinematics the two decay planes coincide to form a single decay plane. So for back-to-back kinematics $\phi=0$ Kim:2023ohr . No integration over $\phi$ is involved to arrive at the final observables in back-to-back kinematics.
@@
Out of these, the first and second conditions are primary ones, while the rest arise as a consequence of the first two criteria. However, none of these quantities are physically observable. The first two conditions also imply that $E_{+}=E_{-}=E_{\mu}\equiv m_{B}/2-E_{\nu}$ which is experimentally observable. Any error, on the muon energy measurement would imply that the muon energy distribution as shown in Fig. 5(c) of Ref. Kim:2021dyj would involve energy bins with bin width corresponding to the experimental error within which equality of $E_{+}$ and $E_{-}$ is satisfied. There would also possibly be extremely slight deviation from the $180^{\circ}$ angle between the two final muons, in an experimental back-to-back realization. This can lead to small deviation of $\Theta$ from $\pi$, say $\Theta=\pi\pm\Delta\Theta$. This implies $\cos\Theta=-\cos\Delta\Theta\approx-1+\Delta\Theta^{2}/2$, so that the error in $\cos\Theta$ measurement is $\Delta\Theta^{2}/2$ and it gets multiplied to the back-to-back muon energy distribution of Eq. (50) of Ref. Kim:2021dyj to give the amount of smearing one can expect from the measurement. Thus, a real experimental realization of the back-to-back kinematics would lead to a muon energy distribution curve similar to Fig. 5(c) of Ref. Kim:2021dyj but it would be a histogram plot with bin size determined by the muon energy resolution and there will be some vertical smearing arising from the slight deviation from $180^{\circ}$ angle requirement. However, the difference between Dirac and Majorana neutrinos should not get washed away as a result of such a tiny smearing from experimental measurements.